# Low-curvature quantum corrections from unitary evolution of de Sitter space

### v1

@@ I Introduction
One of the main obstacles in our attempts to find a consistent quantum framework for gravity arises from the different concepts of time in general relativity and quantum mechanics. While general relativity allows general transformations from one coordinate frame to another, quantum mechanics relies on notions of time evolution and unitarity based on a more restricted concept of time. One immediate difficulty arises from the fact that the (bulk) Hamiltonian of a generally covariant theory vanishes due to diffeomorphism invariance, obstructing any interpretation of such a Hamiltonian as generator of time evolution. These issues have long been discussed in the literature as the problem of time [1, 2, 3].
@@
Assuming FLRW symmetry means the geometric degrees of freedom are reduced to just the scale factor, here seen as evolving in unimodular time. The resulting quantum theories are simple and can be defined straightforwardly. We will see how unitary evolution enforces a “resolution” of the de Sitter horizon, replacing it by a highly quantum region that interpolates between two classical solutions each associated to one half of de Sitter spacetime. We will work both in a traditional Wheeler–DeWitt quantisation and in LQC, where many results show resolution of classical singularities through loop quantum gravity effects [21, 22], so one could ask whether that approach also shows “fake singularity resolution”. Specifically, we will follow the proposal of [23] for defining LQC for unimodular gravity. We will find agreement with the general conclusions of that article in that loop quantum corrections are small assuming that the cosmological constant is far away from the Planck scale, so the qualitiative features are similar to those of the Wheeler–DeWitt theory. The fact that the quantum evolution continues through a highly quantum region replacing the classical horizon was however missed in [23] due to some errors in the analysis, as we will clarify later.
@@ II.1 Hamiltonian unimodular gravity with FLRW symmetry
where $\Gamma^{a}_{i}$ is the torsion-free Levi-Civita connection associated to $E$, $K^{a}_{i}$ is the extrinsic curvature 1-form, $e^{b}_{j}$ is a spatial triad such that $q_{ij}=\delta_{ab}\,e^{a}_{i}e^{b}_{j}$, and $\gamma$ is a free parameter known as the Barbero–Immirzi parameter. In the case of flat FLRW cosmology, these fields take the form [22, 21]
@@
where ${}^{o}e_{i}^{a}$ is a fiducial (co-)triad adapted to the metric $h_{ij}$, i.e., $h_{ij}=\delta_{ab}({}^{o}e^{a}_{i})({}^{o}e^{b}_{j})$, ${}^{o}e_{a}^{i}$ is the corresponding inverse triad, and $V_{0}$ is the volume $V_{0}=\int{\rm d}^{3}x\;\sqrt{h}$ of a finite “fiducial cell” one wants to study. Notice that compared to the metric formulation there is now a new symmetry under the orientation reversal of the fiducial triad.
@@
The final consequence of homogeneity and isotropy is that spatial diffeomorphisms are trivially satisfied, and the Hamiltonian reduces to one constraint that generates time reparametrisations:
@@
In the interest of choosing $T$ as the internal clock, in the following we will fix the lapse to $N=\frac{4}{\kappa\gamma\hbar|\nu|}$. Thus, we obtain the classical Hamiltonian
@@
which is constrained to vanish. (With this lapse choice, indeed $\dot{T}=\{T,H\}=1$.)
@@ II.2 Foliations of de Sitter spacetime
In cosmology it is more common to adopt cosmic time $t$ (choosing $N=1$) and write the FLRW line element in polar spatial coordinates as
@@
$D$-dimensional de Sitter space is a spacetime of maximal symmetry, so it has additional isometries beyond the ones of general FLRW spacetimes. There are then different representations of de Sitter space as FLRW geometries, corresponding to different foliations by spatial slices of constant curvature. In particular, there is a global slicing by positively curved ($k=1$) slices which covers the entire spacetime, and another slicing by flat ($k=0$) slices which covers only half of the spacetime. Many more details on various aspects of de Sitter spacetime can be found, e.g., in [29].
$D$-dimensional de Sitter space can be embedded into a flat $D+1$-dimensional Minkowski spacetime with metric
${\rm d}s_{\textrm{embed}}^{2}=-{\rm d}X_{0}^{2}+{\rm d}X_{1}^{2}+\ldots+{\rm d}X_{D}^{2}\,.$ (17)
In this context de Sitter space is the timelike hyperboloid defined by
$X_{\mu}X^{\mu}=\ell^{2}$ (18)
with $\ell^{2}=3/\Lambda$. This is represented for $D=2$ in figure 1(a). The global coordinates
@@
where $z_{i}z^{i}=1$, cover the entire hyperbola. With these coordinates the metric is described through the FLRW line element (16) with $a(t)=\ell\cosh(t/\ell)$ and $k=1$, and so this is a closed slicing. On the other hand, in the context of cosmology it is most usual to adopt a flat slicing, as we did in the previous section, which can be described through (16) when $k=0$ and $a(t)=e^{t/\ell}$. (To compare with the discussion below (15), recall that $a\sim|\nu|^{1/3}$.) This is accomplished through the choice of coordinates
@@ III Wheeler–DeWitt quantisation
$\Psi(\nu,T)=N\frac{e^{-\frac{\Lambda_{0}}{2\sigma^{2}}}}{2f}\left\{4+\sqrt{\frac{\pi}{f}}\left(ge^{\frac{g^{2}}{4f}}\left[1+\text{Erf}\left(\frac{g}{2\sqrt{f}}\right)\right]+\bar{g}e^{\frac{\bar{g}^{2}}{4f}}\left[1+\text{Erf}\left(\frac{\bar{g}}{2\sqrt{f}}\right)\right]\right)\right\}\,.$ (33)
@@
One can then perform the integral (26) numerically. Choosing some example values for the parameters of the Gaussian, we compute this expectation value and plot it alongside the classical trajectories in Figure 2. As expected and alluded to in the introduction, the quantum trajectory deviates from the classical one close to the de Sitter horizon, where it interpolates between the two classical solutions corresponding to different halves of the whole de Sitter spacetime. This is a consequence of imposing unitarity in a region where the classical solution goes to zero.
@@
This situation is analogous to computing the expectation value of the absolute value of the position of a wavepacket for a free particle moving in one direction. The position variable $x$ can take on positive and negative values, as can $\nu$ in our case. For a wavepacket, $\langle\hat{x}\rangle$ would follow the classical trajectory and eventually reach and cross $x=0$. However, once we consider $|\hat{x}|$, the negative and positive $x$ parts of the wavefunction both contribute positively to $\langle|\hat{x}|\rangle$, and $\langle|\hat{x}|\rangle$ can never reach exactly 0. The trajectory of $\langle|\hat{x}|\rangle$ therefore closely tracks the classical trajectory of $|x|$ when far from $x=0$, but as it approaches this point (of no special significance in the classical trajectory), it necessarily deviates from it. Unitarity does not allow for the quantum trajectory to simply end there and $\langle|\hat{x}|\rangle$ “bounces” close to $x=0$, resulting in the effect we observe in Figure 2. Here, because of the symmetry $\Psi(\nu,T)=\Psi(-\nu,T)$, our state is the equivalent of a superposition of two wavepackets that are mirror images of each other, moving in opposite directions. Both of them will contribute to $\langle|\hat{x}|\rangle$ equally.
While this is all clear from the analogy with a free particle and not a surprise from the quantum point of view, the suggested implications for quantum gravity seem dramatic. The result suggests that the classical horizon (which is foliation-dependent with nothing special happening locally) is replaced by a quantum region that deviates strongly from the classical solution: the ratio of quantum and classical volumes $\langle|\hat{\nu}(T)|\rangle/|\nu_{{\rm cl}}(T)|$, where $\nu_{{\rm cl}}(T)$ is the classical solution corresponding to our chosen parameter values, evidently diverges as $T\rightarrow 0$. This conclusion seems inevitable for any choice of semiclassical state, regardless of how sharply peaked it was chosen to be away from $T=0$; the qualitative features seen in Figure 2 are not sensitive to the details in the state.
In contrast, the analysis of [23, Sec. VI] suggests that the expectation value $\langle|\hat{\nu}|\rangle$, and hence the volume of spatial slices, always remains linear in $T$ following the classical solution (15) for positive $T$ (but then turning negative, which is impossible). Closer inspection of that paper (e.g., Eq. (6.7)) shows that the authors actually calculate the expectation value of $\hat{\nu}$ rather than of $|\hat{\nu}|$, which would be zero at all times due to the assumed symmetry.
@@
Figure 2: Expectation value of $|\hat{\nu}|$ as a function of $T$ for Wheeler–DeWitt theory and LQC, calculated numerically for a state characterised by a Gaussian profile (29) with $\Lambda_{0}=1$ and $\sigma=0.1$, along with dispersions and corresponding classical trajectories. Here $\hbar=G=1$ so that $\kappa=8\pi$; we also set $\lambda=\sqrt{4\sqrt{3}\pi\gamma}$ and $\gamma=0.2375$ as discussed below (37). Then $\Lambda_{p}\approx 10.29$.

### v2

@@ I Introduction
One of the main obstacles in our attempts to find a consistent quantum framework for gravity arises from the different concepts of time in general relativity and quantum mechanics. While general relativity allows general transformations from one coordinate frame to another, quantum mechanics relies on notions of time evolution and unitarity based on a more restricted concept of time. One immediate difficulty arises from the fact that the (bulk) Hamiltonian of a generally covariant theory vanishes due to diffeomorphism invariance, obstructing any interpretation of such a Hamiltonian as generator of time evolution. These issues have long been discussed in the literature as the problem of time [1, 2, 3, 4, 5].
@@
Assuming FLRW symmetry means the geometric degrees of freedom are reduced to just the scale factor, here seen as evolving in unimodular time. The resulting quantum theories are simple and can be defined straightforwardly. We will see how unitary evolution enforces a “resolution” of the de Sitter horizon, replacing it by a highly quantum region that interpolates between two classical solutions each associated to one half of de Sitter spacetime. We will work both in a traditional Wheeler–DeWitt quantisation and in LQC, where many results show resolution of classical singularities through loop quantum gravity effects [23, 24, 25], so one could ask whether that approach also shows “fake singularity resolution”. Specifically, we will follow the proposal of [26] for defining LQC for unimodular gravity. We will find agreement with the general conclusions of that article in that loop quantum corrections are small assuming that the cosmological constant is far away from the Planck scale, so the qualitiative features are similar to those of the Wheeler–DeWitt theory. The fact that the quantum evolution continues through a highly quantum region replacing the classical horizon was however missed in [26] due to some errors in the analysis, as we will clarify later.
In section II we summarise the classical formalism of unimodular gravity for FLRW geometries and written in Ashtekar–Barbero variables. We also review the flat and closed slicings of de Sitter spacetime. Section III discusses the Wheeler–DeWitt quantisation of the model, a choice of Gaussian state and derivation of expectation values and standard deviations in the volume. We show our main result, namely strong quantum effects at the classical horizon. Section IV extends the results to LQC, where we encounter some new features such as a maximum value for the cosmological constant and a quantum correction to $\Lambda$ coming from fundamental discreteness. Qualitative features of our solutions are similar to the Wheeler–DeWitt case. We conclude in section V. Appendix A contains some technical details of the LQC calculations.
@@ II.1 Hamiltonian unimodular gravity with FLRW symmetry
where $\Gamma^{a}_{i}$ is the torsion-free Levi-Civita connection associated to $E$, $K^{a}_{i}$ is the extrinsic curvature 1-form, $e^{b}_{j}$ is a spatial triad such that $q_{ij}=\delta_{ab}\,e^{a}_{i}e^{b}_{j}$, and $\gamma$ is a free parameter known as the Barbero–Immirzi parameter. In the case of flat FLRW cosmology, these fields take the form
@@
where ${}^{o}e_{i}^{a}$ is a fiducial (co-)triad adapted to the metric $h_{ij}$, i.e., $h_{ij}=\delta_{ab}({}^{o}e^{a}_{i})({}^{o}e^{b}_{j})$, and ${}^{o}e_{a}^{i}$ is the corresponding inverse triad. Notice that compared to the metric formulation there is now a new symmetry under the orientation reversal of the fiducial triad. The quantity $V_{0}$ corresponds to the coordinate volume $V_{0}=\int{\rm d}^{3}x\;\sqrt{h}$ of a finite “fiducial cell” one wants to study. This cell could be the entire spatial slice if one assumes compact topology, such as a 3-torus, or a compact subset in case of infinite spatial slices. In the case of non-compact topology, the introduction of a fiducial cell is required to obtain regular expressions since otherwise the total action or Hamiltonian would diverge. The powers of $V_{0}$ appearing in (9) are needed to ensure that none of the properties of the resulting Hamiltonian system depend on the value of $V_{0}$, which is coordinate dependent. More details on these technical aspects and the general construction can be found in [24, 23].
@@
The final consequence of homogeneity and isotropy is that spatial diffeomorphisms are trivially satisfied, and the Hamiltonian reduces to the product of the lapse $N$ with a function $C$ given by
@@
The function $C$ is constrained to vanish through the equation $\partial H/\partial N=0$, and it is a generator of time reparametrisations on the constraint hypersurface in phase space. In the interest of choosing $T$ as the internal clock, in the following we will fix the lapse to $N=\frac{4}{\kappa\gamma\hbar|\nu|}$. Thus, we obtain the classical Hamiltonian
@@
which is again constrained to vanish. (With this lapse choice, indeed $\dot{T}=\{T,H\}=1$.)
@@ II.2 Foliations of de Sitter spacetime
Let us briefly review the different foliations of de Sitter spacetime; this is standard material that can be found in [32], for example. In cosmology it is more common to adopt cosmic time $t$ (choosing $N=1$) and write the FLRW line element in polar spatial coordinates as
@@
$D$-dimensional de Sitter space is a spacetime of maximal symmetry, so it has additional isometries beyond the ones of general FLRW spacetimes. There are then different representations of de Sitter space as FLRW geometries, corresponding to different foliations by spatial slices of constant curvature. In particular, there is a global slicing by positively curved ($k=1$) slices which covers the entire spacetime, and another slicing by flat ($k=0$) slices which covers only half of the spacetime. Many more details on various aspects of de Sitter spacetime can also be found, e.g., in [33].
$D$-dimensional de Sitter space can be embedded into a flat $D+1$-dimensional Minkowski spacetime, where it is defined as the timelike hyperboloid of points satisfying $X_{\mu}X^{\mu}=\ell^{2}$ with $\mu=0,\ldots,D$ and $\ell^{2}=3/\Lambda$. This is represented for $D=2$ in figure 1(a). The global coordinates
@@
where $z_{i}z^{i}=1$, cover the entire hyperbola. With these coordinates the metric is described through the FLRW line element (16) with $a(t)=\ell\cosh(t/\ell)$ and $k=1$, and so this is a closed slicing. On the other hand, there is also the more common flat slicing used here, which can be described through (16) when $k=0$ and $a(t)=e^{t/\ell}$. (To compare with the discussion below (15), recall that $a\sim|\nu|^{1/3}$.) This is accomplished through the choice of coordinates
@@ III Wheeler–DeWitt quantisation
$\Psi(\nu,T)=N\frac{e^{-\frac{\Lambda_{0}}{2\sigma^{2}}}}{2f}\left\{4+\sqrt{\frac{\pi}{f}}\left(ge^{\frac{g^{2}}{4f}}\left[1+\text{Erf}\left(\frac{g}{2\sqrt{f}}\right)\right]+\bar{g}e^{\frac{\bar{g}^{2}}{4f}}\left[1+\text{Erf}\left(\frac{\bar{g}}{2\sqrt{f}}\right)\right]\right)\right\}$ (31)
where $\text{Erf}(x)$ denotes the error function.
@@
One can then perform the integral (24) numerically. Choosing some example values for the parameters of the Gaussian, we compute this expectation value and plot it alongside the classical trajectories in figure 2. We also include error bars showing the standard deviation in $|\nu|$, $\Delta|\nu|=\sqrt{\langle\widehat{|\nu|^{2}}\rangle-\langle\widehat{|\nu|}\rangle^{2}}$. As expected and alluded to in the introduction, the quantum trajectory deviates from the classical one close to the de Sitter horizon, where it interpolates between the two classical solutions corresponding to different halves of the whole de Sitter spacetime. This is a consequence of imposing unitarity in a region where the classical solution goes to zero.
@@
This situation is analogous to computing the expectation value of the absolute value of the position of a wavepacket for a free particle moving in one direction. The position variable $x$ can take on positive and negative values, as can $\nu$ in our case. For a wavepacket, $\langle\hat{x}\rangle$ would follow the classical trajectory and eventually reach and cross $x=0$. However, once we consider $|\hat{x}|$, the negative and positive $x$ parts of the wavefunction both contribute positively to $\langle|\hat{x}|\rangle$, and $\langle|\hat{x}|\rangle$ can never reach exactly 0. The trajectory of $\langle|\hat{x}|\rangle$ therefore closely tracks the classical trajectory of $|x|$ when far from $x=0$, but as it approaches this point (of no special significance in the classical trajectory), it necessarily deviates from it. Unitarity does not allow for the quantum trajectory to simply end there and $\langle|\hat{x}|\rangle$ “bounces” close to $x=0$, resulting in the effect we observe in figure 2. Here, because of the symmetry $\Psi(\nu,T)=\Psi(-\nu,T)$, our state contains two wavepackets corresponding to an expanding and a contracting de Sitter solution; this is the equivalent of a superposition of two wavepackets that are mirror images of each other, moving in opposite directions. Both of them will contribute to $\langle|\hat{x}|\rangle$ equally.
@@
Figure 2: Expectation value of $|\hat{\nu}|$ as a function of $T$ for Wheeler–DeWitt theory and LQC for a state characterised by a Gaussian profile (27) with $\Lambda_{0}=1$ and $\sigma=0.1$, shown with the standard deviation and corresponding classical trajectories. Here $\hbar=G=1$ so that $\kappa=8\pi$; we also set $\lambda=\sqrt{4\sqrt{3}\pi\gamma}$ and $\gamma=0.2375$ as discussed below (36). Then $\Lambda_{p}\approx 10.29$.
While this is all clear from the analogy with a free particle and not a surprise from the quantum point of view, the suggested implications for quantum gravity seem dramatic. The result suggests that the classical horizon (which is foliation-dependent with nothing special happening locally) is replaced by a quantum region that deviates strongly from the classical solution: the ratio of quantum and classical volumes $\langle|\hat{\nu}(T)|\rangle/|\nu_{{\rm cl}}(T)|$, where $\nu_{{\rm cl}}(T)$ is the classical solution corresponding to our chosen parameter values, evidently diverges as $T\rightarrow 0$. This conclusion seems inevitable for any choice of semiclassical state, regardless of how sharply peaked it was chosen to be away from $T=0$; the qualitative features seen in figure 2 are not sensitive to the details in the state, as long as it is semiclassical. Away from $T=0$, expectation values closely follow the classical solution, while there are strong departures near $T=0$.
More quantitatively, we can compute $\langle|\hat{\nu}(0)|\rangle$ using the approximation (32). For $\sigma\ll\sqrt{\Lambda_{0}}$,
$\langle|\hat{\nu}(0)|\rangle\approx\sqrt{\frac{3}{\pi}}\,\frac{2}{\gamma\sigma}\,.$ (33)
The scaling of the minimal value of the volume in the quantum theory with $1/\sigma$ agrees with that seen in similar constructions in quantum cosmology such as [36]. In particular, the precise value depends on the details of the state. However, given that the classical volume goes to zero here, we would see any non-zero value as a strong quantum departure from the classical solution. The dependence on $\sigma$ is more explicitly illustrated in figure 3. We have checked in particular that the results for $\langle|\hat{\nu}(0)|\rangle$ match with the analytical approximation (33).
Figure 3: Expectation value of $|\hat{\nu}|$ as a function of $T$ for Wheeler–DeWitt theory for a state with Gaussian profile (27) with different values for $\sigma$ and $\Lambda_{0}=1$, plotted along with standard deviations and the corresponding classical trajectory with $\Lambda=\Lambda_{0}$ (black lines).
As a last point, we notice that the dispersion in $|\nu|$ grows away from $T=0$, as indicated by an increasing length of the error bars shown in the previously discussed figures. However, if one looks at the relative standard deviation $\Delta|\nu|/\langle|\hat{\nu}|\rangle$, a dimensionless ratio which might be more meaningful to study, one can see that it is actually greatest around the quantum bounce and then falls off as the volume grows towards the future or past. This is also illustrated in figure 4.
Figure 4: Relative standard deviation $\Delta|\nu|/\langle|\hat{\nu}|\rangle$ as a function of $T$ for Wheeler–DeWitt theory and LQC, with state parameters chosen as in figure 2.
In contrast to the results presented here, the analysis of [26, Sec. VI] suggests that the expectation value $\langle|\hat{\nu}|\rangle$, and hence the volume of spatial slices, always remains linear in $T$ following the classical solution (15) for positive $T$, so it does eventually go to zero (and then turns negative, which is impossible). Closer inspection of that paper (e.g., Eq. (6.7)) shows that the authors actually calculate the expectation value of $\hat{\nu}$ rather than of $|\hat{\nu}|$, which would be zero at all times due to the assumed symmetry.